While on a train moving east at a speed of 29.0 m/s relative to the tracks, a passenger walks to the back of the train at a speed of 1.60 m/s relative to the train. Relative to the ground, how far east does the passenger travel in 52.0 seconds? Calculate the answer in meters (m) and rounded to three significant figures.
Answer #1

Speed of passenger relative to the ground = 29 - 1.6 = 27.4 m/s to east

Distance travelled by passenger relating to the ground = 27.4*52 = 1424.8 m = 1420 m ( in 3 significant figures)